# MAB-Based Channel Scheduling for Asynchronous Federated Learning in Non-Stationary Environments

### v1

@@ V ADAPTIVE channel matching BASED ON MARGINAL CONTRIBUTION AND FAIRNESS
Under non-stationary channel conditions, the aggregation state of the federated model becomes inherently unpredictable. From the perspective of performance fairness, the federated model may exhibit a bias toward clients with frequent access to aggregation, thereby degrading the performance of clients assigned to unfavorable channels. Moreover, in terms of collaborative fairness, clients with non-iid distributed data characteristics demonstrate varying marginal contributions, and non-stationary channel states may prevent critical clients from participating in the latest global aggregation. To mitigate these challenges, we propose an adaptive channel allocation strategy that incorporates both marginal contribution and collaborative fairness considerations.
Considering a scenario with $N$ sub-channels and $M$ clients, our approach involves scheduling $M$ sub-channels after performing MAB-based channel scheduling in the $t$-th round of aggregation. In piecewise-stationary scenario, we sort each sub-channel according to the UCB value calculated based on Eq. (26). i.e.
@@
For extremely non-stationary channels, in the M-exp3 algorithm we schedule according to the super-arm, i.e., the combination of scheduling channels in each round. Given the absence of independent information specific to individual channels in each round, we rank the $M$ sub-channels according to their historical mean values. i.e.
@@
where $S$ denotes the overall client set and $S_{-m}$ denotes the set that excluding the client $m$. $U(\cdot)$ is the utility function that measuring the benefit of the client set on global convergence. Due to the computational complexity of directly calculating Shapley values, Jiang et al.[32] propose an estimation terms of the marginal contribution incorporating cosine similarity and model errorr. i.e.
@@
This prioritization facilitates adaptive tuning of channel allocation. When the disparity in staleness among clients is low, indicating relatively equal participation in global aggregation, channel allocation prioritizes efficiency. In this scenario, clients with greater marginal benefits can access better channels. Conversely, when the variance of AoI is significant, indicating some clients are notably lagging, channel allocation shifts toward balancing staleness and enhancing fairness. Thus, clients with higher AoI are afforded the opportunity to access superior channels. The overall procedure for our proposed scheduling strategy is outlined in Fig.1.
@@ VI-A Experimental Setup
To simulate non-IID data distributions, we employ the Dirichlet distribution following [36]. Specifically, we sample $p_{k}\sim\text{Dir}N(\alpha)$ and allocate a proportion $p{k,j}$ of class $k$ instances to client $j$, where $\text{Dir}(\cdot)$ denotes the Dirichlet distribution and $\alpha$ is a concentration parameter ($\alpha>0$). The parameter $\alpha$ controls the degree of data heterogeneity: as $\alpha\to\infty$, the data distribution approaches IID, while $\alpha\to 0$ results in extreme Non-IID conditions.
The experimental setup considers an anarchic and non-stationary channel environment where the central server manages multiple sub-channels without prior knowledge of the time-varying CSI. The channels are vulnerable to potential attacks, leading to state mutations during communication rounds with $C_{T}$ breakpoints over $T$ rounds. We evaluate our algorithm under two distinct non-stationary scenarios: an extremely dynamic scenario where channel states may change abruptly at any round without warning, and a piece-wise stationary scenario where channel characteristics maintain statistical stationarity between breakpoints. For comparative analysis, we implement a random scheduling baseline where the central server performs random channel matching and client-channel matching in each communication round.
@@ VI-B Performance Comparisons
As demonstrated in Fig.2(a), GLR-CUCB and M-exp3 algorithms exhibit superior performance over random scheduling in identifying high-quality channels and minimizing AoI regret. Both algorithms achieve sub-linear growth in AoI regret, aligning with our theoretical analysis. The integration of AoI awareness enables dynamic channel scheduling based on AoI information, resulting in further reduction of AoI regret. Additionally, GLR-CUCB demonstrates better AoI regret performance compared to M-exp3, consistent with Theorems 2 and 3.
@@
Performance of Test accuracy: We set $\alpha=0.5$ to control the degree of non-iid between clients. To assess the effectiveness of our proposed approach, we conducted ablation experiments examining both the channel scheduling strategy and adaptive channel matching on federated model convergence. The experiments were conducted under two different scales. For piece-wise stationary channel environments, we set $N=30$ channels and $M=20$ clients. Based on our regret analysis results showing the impact of system scale on M-exp3 algorithm performance, we evaluated the extremely non-stationary channel environments in a smaller-scale communication system with $N=6$ channels and $M=4$ clients.
@@
As shown in Fig. 3, experimental results on AoI regret validate that GLR-CUCB and M-exp3 enhance communication efficiency and accelerate federated model convergence by mitigating AoI, consistent with our theoretical analysis. In piecewise stationary channels, GLR-CUCB with aware allocation achieves approximately 55% accuracy on CIFAR-10 and 39% on CIFAR-100, outperforming random scheduling. Moreover, our proposed algorithms exhibit faster convergence, reaching stable accuracy within 100 communication rounds, compared to 150 rounds for random scheduling. In extremely non-stationary channels, M-exp3 with aware allocation maintains a stable performance of approximately 58% accuracy on CIFAR-10 and 38% on CIFAR-100, demonstrating improvements over random scheduling and achieving convergence within 75 rounds instead of 100.
@@
The ablation experiments further demonstrate the effectiveness of adaptive channel matching. Comparing the performance between random allocation and aware allocation, we observe consistent improvements across all scenarios. Specifically, in piece-wise stationary channels, aware allocation provides an additional accuracy gain for CIFAR-10 and for CIFAR-100 compared to random allocation. This enhancement in performance indicates that the proposed algorithm effectively balances the staleness among clients and mitigates the issue of clients falling significantly behind the global process.
@@
Performance of fairness: As evidenced by the AoI variance results in Figures 4, our proposed algorithms significantly reduce the accumulated AoI variance. In piece-wise stationary channels, GLR-CUCB with aware allocation reduces the cumulative AoI variance by approximately 70% (from 450 to 150) on CIFAR-10 and 75% (from 350 to 90) on CIFAR-100 compared to random scheduling at round 250. Similarly, in extremely non-stationary channels, M-exp3 with aware allocation achieves a reduction of about 75% (from 400 to 100) on CIFAR-10 and 70% (from 800 to 240) on CIFAR-100.
@@
From the experimental results, it can be seen that the cumulative AoI variance of the proposed algorithms tends to flatten out with the training process, i.e., the difference in the AoI of the individual clients nearly ceases to change. Specifically, after around 150 communication rounds, the slope of AoI variance curve with aware allocation becomes significantly smaller (approximately 0.2) compared to random scheduling (approximately 1.5). We find that the GLR-CUCB as well as M-exp3 algorithms also reduce the AoI variance to some extent, which mainly stems from the ability of the algorithms to reduce the overall AoI of the clients, achieving a 40-50% reduction compared to random scheduling. When certain clients have significantly lagged behind the global update, the adaptive algorithms can dynamically adjust the priority coefficients of clients, allowing the lagging clients to be prioritized for better channels, thereby mitigating client drift. The effectiveness of this approach is particularly evident in the extremely non-stationary scenario, where aware allocation maintains the cumulative AoI variance below 100 for CIFAR-10 and 250 for CIFAR-100, compared to random scheduling which reaches 400 and 800 respectively. By balancing the aggregation participation of individual clients, the algorithms facilitate a more fair aggregation process.
@@ VII Conclusions
In non-stationary wireless environments, asynchronous federated learning faces delayed updates and client drift challenges. We propose a MAB-based channel scheduling framework with theoretical analysis for two non-stationary scenarios, establishing AoI regret bounds for GLR-CUCB and M-exp3 algorithms. To address imbalanced client updates, we introduce an adaptive channel matching strategy considering marginal utility and fairness. Empirical results demonstrate enhanced communication efficiency with maintained model convergence and fair client participation in non-stationary federated learning.
@@ Appendix A Proof of Theorem 1
$\hskip 14.22636pt+\underbrace{\frac{L\eta^{2}\lambda^{2}}{2}\mathbb{E}\left\|\frac{1}{|S_{t}|\lambda}\sum_{k\in S_{t}}\sum_{l=0}^{\lambda-1}\tilde{\nabla}F_{k}\left(\boldsymbol{w}_{k,t-a_{i}(t)}^{(l)}\right)\right\|^{2}}_{A_{2}}$ (44)
where $o_{t}=\frac{1}{|S_{t}|\lambda}\sum_{k\in S_{t}}\sum_{l=0}^{\lambda-1}\tilde{\nabla}F_{k}\left(\boldsymbol{w}_{k,t-a_{i}(t)}^{(l)}\right)-\frac{1}{M\lambda}\sum_{i=1}^{M}\sum_{l=0}^{\lambda-1}\tilde{\nabla}F_{k}\left(\boldsymbol{w}_{k,t-a_{i}(t)}^{(l)}\right)$. In the following, we bound the terms $A_{1}$ and $A_{2}$ respectively.
@@
$+\frac{3}{M}\sum_{i=1}^{M}L^{2}\mathbb{E}\left\|\boldsymbol{w}_{t-a_{i}(t)}-\boldsymbol{w}_{t}\right\|^{2}+3\left\|\nabla F\left(\boldsymbol{w}_{t}\right)\right\|^{2}.$ (51)
@@
Similarly, we can bound $A_{1}$ as follows
@@
$\begin{array}[]{l}\mathbb{E}\left\|\boldsymbol{w}_{m,t}-\boldsymbol{w}_{m,t^{\prime}}\right\|^{2}=\mathbb{E}\left\|\sum_{j=t^{\prime}}^{t-1}\left(\boldsymbol{w}_{j+1}-\boldsymbol{w}_{j}\right)\right\|^{2}\\ =\tilde{\eta}^{2}\mathbb{E}\left\|\sum_{j=t^{\prime}}^{t-1}\frac{1}{\lambda|S_{j}|}\sum_{k\in S_{j}}\sum_{l=0}^{\lambda-1}\tilde{\nabla}F_{k}\left(\boldsymbol{w}_{k,j-\tau_{k,j}}^{(l)}\right)\right\|^{2}\\ \stackrel{{\scriptstyle(a)}}{{\leq}}\tilde{\eta}^{2}\left(t-t^{\prime}\right)\sum\limits_{j=t^{\prime}}^{t-1}\mathbb{E}\left\|\frac{1}{\lambda|S_{j}|}\sum\limits_{k\in S_{j}}\sum\limits_{l=0}^{\lambda-1}\tilde{\nabla}F_{k}\left(\boldsymbol{w}_{k,j-\tau_{k,j}}^{(l)}\right)\right\|^{2}\\ \stackrel{{\scriptstyle(b)}}{{\leq}}2\tilde{\eta}^{2}\left(t-t^{\prime}\right)\sum\limits_{j=t^{\prime}}^{t-1}\mathbb{E}\|\frac{1}{\lambda|S_{j}|}\sum\limits_{k\in S_{j}}\sum\limits_{l=0}^{\lambda-1}\tilde{\nabla}F_{k}\left(\boldsymbol{w}_{k,j-\tau_{k,j}}^{(l)}\right)\\ \hskip 14.22636pt-\nabla F_{k}\left(\boldsymbol{w}_{k,j-\tau_{k,j}}^{(l)}\right)\|^{2}+2\tilde{\eta}^{2}\left(t-t^{\prime}\right)^{2}G^{2}\\ \leq 2\tilde{\eta}^{2}\left(t-t^{\prime}\right)^{2}\left(\sigma^{2}+G^{2}\right)\end{array}$ (54)
@@ Appendix B Proof of Theorem 2
This theorem can be proved by modifying the proof of [37, Theorem 3]. The sequence of Good (1) and Bad (0) channel states can be sampled from the Bernoulli distribution $\mathcal{B}_{\mu}$ where $\mu$ denotes the mean. We consider the arm configuration $\textbf{F}_{(\textbf{i})}$ in which there are only one best super-arm $\textbf{i}\in C(N,M)$. For any channel $k\in$ i, its mean is $\mathcal{B}_{\frac{1}{2}+\epsilon}$. For any channel $k\notin$ i, its mean is $\mathcal{B}_{\frac{1}{2}}$. $C(N,M)$ denotes the set of $M$ arms randomly selected from $N$ arms and $|C(N,M)|$ denotes the size of $C(N,M)$. The AoI regret of every policy is lower bounded by the regret incurred in this bandit problem.
@@
Let $\mathbb{E}_{(\textbf{i})}$ and $\mathbb{P}_{(\textbf{i})}$ denote the expectation and the probability measure with respect to the arm configuration $\textbf{F}_{(\textbf{i})}$. In an alternative coupled system in [38], $\{U{(t)_{t\geq 1}}\}$ is a i.i.d. random variables distributed uniformly in (0,1). Let ${R_{k}}(t)=\mathbb{I}\{U(t)\leq{\mu_{k}}\}$ denote the transmission result on channel $k$ in round $t$, where $\mathbb{I}[\cdot]$ denotes the indicator function. From (14), (59) and [39, (5)], we can get the AoI regret of client $i$ in round $t$.
@@
$=\frac{2\epsilon}{1+2\epsilon}\sum_{k\notin\textbf{i}}[\mathbb{P}_{(\textbf{i})}(\mathbb{I}\{k_{i}(t)=k\}=1)]$ (61)
Therefore, we can get the cumulative AoI of all clients in $t$ rounds.
@@
Next, we upper bound $\mathbb{E}_{(\textbf{i})}[N_{\textbf{i}}]$. The key idea is to use an arm configuration $\textbf{F}_{(0)}$ with all arms following $\mathcal{B}_{\frac{1}{2}}$ as the benchmark. From [37, Theorem 3], we have
@@
$\begin{split}\frac{\sum_{\textbf{i}\in C(N,M)}\mathbb{E}_{(\textbf{i})}[N_{\textbf{i}}]}{|C(N,M)|}\leq M\left(\frac{MT}{N}+\frac{MT}{2}\sqrt{\frac{T}{N}\log\frac{1}{1-4\epsilon^{2}}}\right)\end{split}$ (66)
@@
$\begin{split}R_{\pi}(T)&\geq\frac{2\epsilon}{1+2\epsilon}M\left(T-\frac{MT}{N}-\frac{MT}{2}\sqrt{\frac{T}{N}\log\frac{1}{1-4\epsilon^{2}}}\right)\\ \end{split}$ (67)
@@
Proof: We consider a special piece-wise stationary channel distribution where the best super-arm only changes once. The first $M$ arm is $\mathcal{B}_{\frac{1}{2}+\epsilon}$ where $\frac{1}{2}+\epsilon$ is the means of Bernoulli distribution. The rest arms are $\mathcal{B}_{\frac{1}{2}}$. The best super arm is denoted as f. If there are breakpoints, we divide $T$ into $K$ intervals of the same size $\tau$. We denote the channel configuration $\textbf{F}_{(j)}$ that the latter $M$ arms change from $\mathcal{B}_{\frac{1}{2}}$ to $\mathcal{B}_{\frac{1}{2}+\epsilon}$ in the $j$ interval$(1\leq j\leq K=[\frac{T}{\tau}])$, and the other arms are $\mathcal{B}_{\frac{1}{2}}$ and I denotes the best super arm. Let $\mathbb{E}_{(j)}$ and $\mathbb{P}_{(j)}$ denote the expectation and probability of $\textbf{F}_{(j)}$. We denote the channel configuration $\textbf{F}_{(0)}$ that the channel of each arm is $\mathcal{B}_{\frac{1}{2}}$, that is, the stationary channel. Let $\mathbb{E}_{(0)}$ and $\mathbb{P}_{(0)}$ denote the expectation and probability of $\textbf{F}_{(0)}$.
@@
$\begin{split}R_{\pi}(T)\geq\frac{2\epsilon}{1+2\epsilon}\left(M\tau-\frac{M\tau}{K}-\frac{M\tau}{2K}\sqrt{M\tau\log\frac{1}{1-4\epsilon^{2}}}\right)\end{split}$ (73)
@@
$\begin{split}R_{\pi}(T)\geq\frac{2\epsilon}{1+2\epsilon}\left(\frac{MT}{N}-\frac{MT}{N^{2}}-\frac{MT}{2N^{2}}\sqrt{\frac{MT}{N}\log\frac{1}{1-4\epsilon^{2}}}\right)\end{split}$ (74)

### v2

@@ V ADAPTIVE CHANNEL MATCHING BASED ON MARGINAL CONTRIBUTION AND FAIRNESS
Under non-stationary channel conditions, the aggregation state of the federated model becomes inherently unpredictable. From the perspective of performance fairness, the federated model may exhibit a bias toward clients with frequent access to aggregation, thereby degrading the performance of clients assigned to unfavorable channels. Moreover, in terms of collaborative fairness, clients with non-iid distributed data characteristics demonstrate varying marginal contributions, and non-stationary channel states may prevent critical clients from participating in the latest global aggregation. To mitigate these challenges, we propose an adaptive channel matching strategy that incorporates both marginal contribution and collaborative fairness considerations.
Considering a scenario with $N$ sub-channels and $M$ clients, our approach involves scheduling $M$ sub-channels after performing MAB-based channel scheduling in the $t$-th round of aggregation. In piecewise-stationary scenario, we sort each sub-channel according to the UCB value calculated based on Eq. (26), i.e.,
@@
For extremely non-stationary channels, in the M-exp3 algorithm we schedule according to the super-arm, i.e., the combination of scheduling channels in each round. Given the absence of independent information specific to individual channels in each round, we rank the $M$ sub-channels according to their historical mean values, i.e.,
@@
where $S$ denotes the overall client set and $S_{-m}$ denotes the set that excluding the client $m$. $U(\cdot)$ is the utility function that measuring the benefit of the client set on global convergence. Due to the computational complexity of directly calculating Shapley values, an estimation method based on marginal contributions that incorporates cosine similarity and model error has been developed [33]. This leads to the following formulation:
@@
This prioritization facilitates adaptive tuning of channel matching. When the disparity in staleness among clients is low, indicating relatively equal participation in global aggregation, channel matching prioritizes efficiency. In this scenario, clients with greater marginal benefits can access better channels. Conversely, when the variance of AoI is significant, indicating some clients are notably lagging, channel matching shifts toward balancing staleness and enhancing fairness. Thus, clients with higher AoI are afforded the opportunity to access superior channels. The overall procedure for our proposed scheduling strategy is outlined in Fig.1.
@@ VI-A Experimental Setup
To simulate non-IID data distributions, we employ the Dirichlet distribution following [38]. Specifically, we sample $p_{k}\sim\text{Dir}N(\alpha)$ and allocate a proportion $p_{k,j}$ of class $k$ instances to client $j$, where $\text{Dir}(\cdot)$ denotes the Dirichlet distribution and $\alpha$ is a concentration parameter ($\alpha>0$). The parameter $\alpha$ controls the degree of data heterogeneity: as $\alpha\to\infty$, the data distribution approaches IID, and $\alpha\to 0$ results in extreme Non-IID conditions.
The experimental setup considers an anarchic and non-stationary channel environment where the central server manages multiple sub-channels without prior knowledge of the time-varying CSI. The channels are vulnerable to potential attacks, leading to state mutations during communication rounds with $C_{T}$ breakpoints over $T$ rounds. We evaluate our algorithm under two distinct non-stationary scenarios: an extremely dynamic scenario where channel states may change abruptly at any round without warning, and a piecewise-stationary scenario where channel characteristics maintain statistical stationarity between breakpoints. For comparative analysis, we implement a random scheduling baseline where the central server performs random channel scheduling and client-channel matching in each communication round.
@@ VI-B Performance Comparisons
As demonstrated in Fig.2(a), GLR-CUCB and M-exp3 algorithms exhibit superior performance over random scheduling in identifying high-quality channels and minimizing AoI regret. Both algorithms achieve sub-linear growth in AoI regret, aligning with our theoretical analysis. The integration of AoI awareness enables dynamic channel scheduling based on AoI information, resulting in further reduction of AoI regret. Additionally, GLR-CUCB demonstrates better AoI regret performance compared to M-exp3, consistent with Theorems 3 and 5.
@@
Performance of Test accuracy: We set $\alpha=0.5$ to control the degree of non-iid between clients. To assess the effectiveness of our proposed approach, we conducted ablation experiments examining both the channel scheduling strategy and adaptive channel matching on federated model convergence. The experiments were conducted under two different scales. For piecewise-stationary channel environments, we set $N=30$ channels and $M=20$ clients. Based on our regret analysis results showing the impact of system scale on M-exp3 algorithm performance, we evaluated the extremely non-stationary channel environments in a smaller-scale communication system with $N=6$ channels and $M=4$ clients.
@@
As shown in Fig. 3, experimental results on AoI regret validate that GLR-CUCB and M-exp3 enhance communication efficiency and accelerate federated model convergence by mitigating AoI, consistent with our theoretical analysis. In piecewise-stationary channels, GLR-CUCB with aware matching achieves approximately 55% accuracy on CIFAR-10 and 39% on CIFAR-100, outperforming random scheduling. Moreover, our proposed algorithms exhibit faster convergence, reaching stable accuracy within 100 communication rounds, compared to 150 rounds for random scheduling. In extremely non-stationary channels, M-exp3 with aware matching maintains a stable performance of approximately 58% accuracy on CIFAR-10 and 38% on CIFAR-100, demonstrating improvements over random scheduling and achieving convergence within 75 rounds instead of 100.
@@
The ablation experiments further demonstrate the effectiveness of adaptive channel matching. Comparing the performance between random matching and aware matching, we observe consistent improvements across all scenarios. Specifically, in piecewise-stationary channels, aware matching provides an additional accuracy gain for CIFAR-10 and for CIFAR-100 compared to random matching. This enhancement in performance indicates that the proposed algorithm effectively balances the staleness among clients and mitigates the issue of clients falling significantly behind the global process.
@@
Performance of fairness: As evidenced by the AoI variance results in Figures 4, our proposed algorithms significantly reduce the accumulated AoI variance. In piecewise-stationary channels, GLR-CUCB with aware matching reduces the cumulative AoI variance by approximately 75% (from 480 to 130) on CIFAR-10 and 77% (from 360 to 80) on CIFAR-100 compared to random scheduling at round 250. Similarly, in extremely non-stationary channels, M-exp3 with aware matching achieves a reduction of about 75% (from 400 to 100) on CIFAR-10 and 71% (from 800 to 230) on CIFAR-100.
@@
From the experimental results, it can be seen that the cumulative AoI variance of the proposed algorithms tends to flatten out with the training process, i.e. the difference in the AoI of the individual clients nearly ceases to change. Specifically, after around 150 communication rounds, the slope of AoI variance curve with aware matching becomes significantly smaller (approximately 0.2) compared to random scheduling (approximately 1.5). We find that the GLR-CUCB as well as M-exp3 algorithms also reduce the AoI variance to some extent, which mainly stems from the ability of the algorithms to reduce the overall AoI of the clients. When certain clients have significantly lagged behind the global update, the adaptive algorithms can dynamically adjust the priority coefficients of clients, allowing the lagging clients to be prioritized for better channels, thereby mitigating client drift. The effectiveness of this approach is particularly evident in the extremely non-stationary scenario, where aware matching maintains the cumulative AoI variance below 100 for CIFAR-10 and 250 for CIFAR-100, compared to random scheduling which reaches 400 and 800 respectively. By balancing the aggregation participation of individual clients, the algorithms facilitate a fairer aggregation process.
@@ VII Conclusions
In non-stationary wireless environments, asynchronous federated learning faces stale updates and client drift challenges. We propose a MAB-based channel scheduling with theoretical analysis for two non-stationary scenarios, establishing AoI regret bounds for GLR-CUCB and M-exp3 algorithms. To address imbalanced client updates, we introduce an adaptive channel matching strategy considering marginal utility and fairness. Empirical results demonstrate enhanced communication efficiency with maintained model convergence and fairer client participation in non-stationary federated learning.
@@ Appendix A Proof of Theorem 1
$\hskip 14.22636pt+\underbrace{\frac{L\eta^{2}\lambda^{2}}{2}\mathbb{E}\left\|\frac{1}{|S_{t}|\lambda}\sum_{k\in S_{t}}\sum_{l=0}^{\lambda-1}\tilde{\nabla}F_{k}\left(\boldsymbol{w}_{k,t-a_{i}(t)}^{(l)}\right)\right\|^{2}}_{A_{2}},$ (44)
where $o_{t}=\frac{1}{|S_{t}|\lambda}\sum_{k\in S_{t}}\sum_{l=0}^{\lambda-1}\tilde{\nabla}F_{k}\left(\boldsymbol{w}_{k,t-a_{i}(t)}^{(l)}\right)-\frac{1}{M\lambda}\sum_{i=1}^{M}\sum_{l=0}^{\lambda-1}\tilde{\nabla}F_{k}\left(\boldsymbol{w}_{k,t-a_{i}(t)}^{(l)}\right)$, ~$\tilde{\nabla}F_{k}(\cdot)$ represents the stochastic gradient and ${\nabla}F_{k}(\cdot)$ denotes the true gradient. In the following, we bound the terms $A_{1}$ and $A_{2}$ respectively.
@@
$+\frac{3}{M}\sum_{i=1}^{M}L^{2}\mathbb{E}\left\|\boldsymbol{w}_{t-a_{i}(t)}-\boldsymbol{w}_{t}\right\|^{2}+3\left\|\nabla F\left(\boldsymbol{w}_{t}\right)\right\|^{2},$ (51)
@@
Similarly, we can bound $A_{1}$ as follows:
@@
$\begin{array}[]{l}\mathbb{E}\left\|\boldsymbol{w}_{m,t}-\boldsymbol{w}_{m,t^{\prime}}\right\|^{2}=\mathbb{E}\left\|\sum_{j=t^{\prime}}^{t-1}\left(\boldsymbol{w}_{j+1}-\boldsymbol{w}_{j}\right)\right\|^{2}\\ =\tilde{\eta}^{2}\mathbb{E}\left\|\sum_{j=t^{\prime}}^{t-1}\frac{1}{\lambda|S_{j}|}\sum_{k\in S_{j}}\sum_{l=0}^{\lambda-1}\tilde{\nabla}F_{k}\left(\boldsymbol{w}_{k,j-\tau_{k,j}}^{(l)}\right)\right\|^{2}\\ {\leq}\tilde{\eta}^{2}\left(t-t^{\prime}\right)\sum\limits_{j=t^{\prime}}^{t-1}\mathbb{E}\left\|\frac{1}{\lambda|S_{j}|}\sum\limits_{k\in S_{j}}\sum\limits_{l=0}^{\lambda-1}\tilde{\nabla}F_{k}\left(\boldsymbol{w}_{k,j-\tau_{k,j}}^{(l)}\right)\right\|^{2}\\ {\leq}2\tilde{\eta}^{2}\left(t-t^{\prime}\right)\sum\limits_{j=t^{\prime}}^{t-1}\mathbb{E}\|\frac{1}{\lambda|S_{j}|}\sum\limits_{k\in S_{j}}\sum\limits_{l=0}^{\lambda-1}\tilde{\nabla}F_{k}\left(\boldsymbol{w}_{k,j-\tau_{k,j}}^{(l)}\right)\\ \hskip 14.22636pt-\nabla F_{k}\left(\boldsymbol{w}_{k,j-\tau_{k,j}}^{(l)}\right)\|^{2}+2\tilde{\eta}^{2}\left(t-t^{\prime}\right)^{2}G^{2}\\ \leq 2\tilde{\eta}^{2}\left(t-t^{\prime}\right)^{2}\left(\sigma^{2}+G^{2}\right)\end{array}$ (54)
@@ Appendix B Proof of Theorem 2
This theorem can be proved by modifying the proof of [35, Theorem 3]. The sequence of Good (1) and Bad (0) channel states can be sampled from the Bernoulli distribution $\mathcal{B}_{\mu}$ where $\mu$ denotes the mean. We consider the arm configuration $\textbf{F}_{(\textbf{i})}$ in which there are only one best super-arm $\textbf{i}\in C(N,M)$. For any channel $k\in$ i, its mean is $\mathcal{B}_{\frac{1}{2}+\epsilon}$. For any channel $k\notin$ i, its mean is $\mathcal{B}_{\frac{1}{2}}$. $C(N,M)$ denotes the set of $M$ arms randomly selected from $N$ arms and $|C(N,M)|$ denotes the size of $C(N,M)$. The AoI regret of every policy is lower bounded by the regret incurred in this bandit problem.
@@
Let $\mathbb{E}_{(\textbf{i})}$ and $\mathbb{P}_{(\textbf{i})}$ denote the expectation and the probability measure with respect to the arm configuration $\textbf{F}_{(\textbf{i})}$. In an alternative coupled system in [39], $\{U{(t)_{t\geq 1}}\}$ is a i.i.d. random variables distributed uniformly in (0,1). Let ${R_{k}}(t)=\mathbb{I}\{U(t)\leq{\mu_{k}}\}$ denote the transmission result on channel $k$ in round $t$, where $\mathbb{I}[\cdot]$ denotes the indicator function. From (14), (59) and [40, (5)], we can get the AoI regret of client $i$ in round $t$.
@@
$=\frac{2\epsilon}{1+2\epsilon}\sum_{k\notin\textbf{i}}[\mathbb{P}_{(\textbf{i})}(\mathbb{I}\{k_{i}(t)=k\}=1)].$ (61)
Therefore, we can get the cumulative AoI of all clients in $t$ rounds:
@@
Next, we upper bound $\mathbb{E}_{(\textbf{i})}[N_{\textbf{i}}]$. The key idea is to use an arm configuration $\textbf{F}_{(0)}$ with all arms following $\mathcal{B}_{\frac{1}{2}}$ as the benchmark. From [35, Theorem 3], we have
@@
$\begin{split}\frac{\sum_{\textbf{i}\in C(N,M)}\mathbb{E}_{(\textbf{i})}[N_{\textbf{i}}]}{|C(N,M)|}\leq M\left(\frac{MT}{N}+\frac{MT}{2}\sqrt{\frac{T}{N}\log\frac{1}{1-4\epsilon^{2}}}\right).\end{split}$ (66)
@@
$\begin{split}R_{\pi}(T)&\geq\frac{2\epsilon}{1+2\epsilon}M\left(T-\frac{MT}{N}-\frac{MT}{2}\sqrt{\frac{T}{N}\log\frac{1}{1-4\epsilon^{2}}}\right).\\ \end{split}$ (67)
@@
Proof: We consider a special piecewise-stationary channel distribution where the best super-arm only changes once. The first $M$ arm is $\mathcal{B}_{\frac{1}{2}+\epsilon}$ where $\frac{1}{2}+\epsilon$ is the means of Bernoulli distribution. The rest arms are $\mathcal{B}_{\frac{1}{2}}$. The best super arm is denoted as f. If there are breakpoints, we divide $T$ into $K$ intervals of the same size $\tau$. We denote the channel configuration $\textbf{F}_{(j)}$ that the latter $M$ arms change from $\mathcal{B}_{\frac{1}{2}}$ to $\mathcal{B}_{\frac{1}{2}+\epsilon}$ in the $j$ interval$(1\leq j\leq K=[\frac{T}{\tau}])$, and the other arms are $\mathcal{B}_{\frac{1}{2}}$ and I denotes the best super arm. Let $\mathbb{E}_{(j)}$ and $\mathbb{P}_{(j)}$ denote the expectation and probability of $\textbf{F}_{(j)}$. We denote the channel configuration $\textbf{F}_{(0)}$ that the channel of each arm is $\mathcal{B}_{\frac{1}{2}}$, that is, the stationary channel. Let $\mathbb{E}_{(0)}$ and $\mathbb{P}_{(0)}$ denote the expectation and probability of $\textbf{F}_{(0)}$.
@@
$\begin{split}R_{\pi}(T)\geq\frac{2\epsilon}{1+2\epsilon}\left(M\tau-\frac{M\tau}{K}-\frac{M\tau}{2K}\sqrt{M\tau\log\frac{1}{1-4\epsilon^{2}}}\right).\end{split}$ (73)
@@
$\begin{split}R_{\pi}(T)\geq\frac{2\epsilon}{1+2\epsilon}\left(\frac{MT}{N}-\frac{MT}{N^{2}}-\frac{MT}{2N^{2}}\sqrt{\frac{MT}{N}\log\frac{1}{1-4\epsilon^{2}}}\right).\end{split}$ (74)